# Ferrotoroidal order of the monolayer ice in martyite

### v1

@@ Abstract
In conclusion, we study the water of crystallization on a quasi 2D honeycomb lattice in martyite, which shows the hierarchical ordering at low temperatures. At room temperature (HT), water molecules are dynamically disordered due to the geometrical frustration of the honeycomb lattice. Here, the rotational and vibrational motions are allowed in the porous framework. At $T_{\mathrm{c1}}$, these water molecules cooperatively form H2O hexamers with hydrogen bonds, leading to the ferrotoroidal order in the LT phase. Once the H2O hexamer is formed, the rotational motions of H2O are forbidden, while the vibrational motions along the hydrogen bonds are still alive. Finally at $T_{\mathrm{c2}}$, these vibrational motions of H2O freeze with the deformation of the H2O hexamers in the LT’ phase. After the deformation, the orientation of the type-1 hexamer (location of the long hydrogen bond) appears as a new degree of freedom. The long hydrogen bonds favour lying close to each other, leading to the formation of the H2O octadecamers. These successive phase transitions gradually break the original honeycomb-lattice symmetry and lead the system close to the ground state, which is difficult to achieve on the highly degenerate energy landscape. This degeneracy is the feature of the monolayer ice, behind the polymorphism of H2O.
@@ Acknowledgements
We acknowledge K. Homma, S. Morita, N. Kawashima, D. Hamane, J. Nasu, and S. Hayami for the fruitful discussions. The synchrotron radiation experiments were performed at SPring-8 with the approval of the Japan Synchrotron Radiation Research Institute (JASRI) (Proposal Nos. 2023A1687 and 2024A1709). The crystal structure figures have been created by using the visualization software VESTA [38].
Funding
This work was partly supported by JSPS KAKENHI, Grant-in-Aid for Scientific Research (Nos. 22K14010, 23H04861, 24K06944, and 24H01644).
Author contributions
T.N. conceived and designed the project. T.N., H.I., T.Y., A.M., and Z.H. grew and characterized the single- and polycrystals of martyite. T.N., M.K., O.Y. and T.M. performed the dielectric measurements. T.N., T.H., and N.O. performed the Raman scattering experiments. S.K., Y.N., and T.A. performed the synchrotron XRD experiments on the single crystal. J.K. and K.K. performed the MD simulations. All authors discussed the results. T.N., S.K., and K.K. wrote the manuscript under the supervision of T.H., Z.H., Y.T., T.A., and T.M.
Competing interests
The authors declare no competing interests.
Data and materials availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.
Experimental
Sample preparation. We obtained single- and polycrystals of martyite by hydrothermal synthesis. We put 100 mg of ZnO (99.0+%, Fujifilm Wako) powder, 200 mg of V2O5 (99.0+%, Fujifilm Wako) powder, 200 mg of citric acid, and 50 ml of deionized water in Teflon-lined stainless steel autoclaves, and kept them at 130 C for 10–30 days. Hexagonal transparent single crystals up to the size of 0.5$\times$0.5$\times$0.1 mm3 were obtained.
@@
Single-crystal XRD. We performed the XRD experiments using a single crystal of 80$\times$70$\times$10 $\mu$m3 on BL02B1 at a synchrotron facility SPring-8 in Japan [39]. We employed a He-gas-blowing device for controlling the temperature between 30 and 300 K. The X-ray wavelength was $\lambda=$ 0.30918 Å. We used a two-dimensional detector CdTe PILATUS to record the diffraction pattern. The intensities of Bragg reflections with the interplane distance $d>0.3$were collected by CrysAlisPro program [40]. Intensities of equivalent reflections were averaged and the structural parameters were refined by using Jana2006 [41].
MD simulation. We carried out molecular dynamics (MD) simulations of two-dimensional water in martyite using GROMACS 2018 software [42]. The model system consists of $N_{w}$ water molecules confined between two atomic layers of martyite. The atomic positions of martyite other than those of water molecules were fixed during the simulations. The number $N_{w}$ of water molecules was 288 unless otherwise stated. The simulation cell was a rectangular prism under periodic boundary conditions in the two dimensions parallel to the atomic layers. The temperature was controlled by the Nosé-Hoover thermostat and the volume of the system was fixed. Other details are described in SM.
Supplemental Material of ”Ferrotoroidal order of the monolayer ice in martyite”
I Crystal growth
We grew single- and polycrystals of martyite by hydrothermal synthesis. We put 100 mg of ZnO (99.0+%, Fujifilm Wako) powder, 200 mg of V2O5 (99.0+%, Fujifilm Wako) powder, 200 mg of citric acid, and 50 ml of deionized water in Teflon-lined stainless steel autoclaves, and kept them at 140 C for 10–30 days. Hexagonal transparent single crystals up to the size of 0.5$\times$0.5$\times$0.1 mm3 were obtained as Fig. S1(a). The edge of the crystal was often red-coloured, probably because of crystallographic defects and water loss. We note that the green part of the photo (Fig. S1(b)) also has a similar crystal structure, although we did not use it for the experiments. The green colour probably indicates that the valence of V ions is shifted from 5+.
Fig. S 1: Photographs of martyite. (a) Single crystals of martyite. (b) Crystal cluster of martyite.
The crystal structure of martyite is topologically the same as the kagomé antiferromagnet, volborthite [Cu3(V2O7)(OH)2$\cdot$2H2O] [43, 44, 45]. The crystal system of volborthite is monoclinic because of the Jahn–Teller distortion of Cu2+ ions. The water of crystal in volborthite is ordered although the molecular orientation changes upon the orbital switching or flipping around 300 K [46]. In contrast, the crystal structure of martyite is trigonal because the $3d$ orbitals of Zn2+ ions are fully occupied and the Jahn–Teller instability is absent. The water of crystal in martyite is located at the perfect honeycomb lattice and is dynamically disordered at room temperature. This difference indicates that the geometrical frustration of the perfect honeycomb lattice suppresses the ordering of H2O molecules.
@@ II Single-crystal XRD
The results of the structural analysis of martyite are summarized in Tables S I–S IV and Figures S2–S5. Figures S2(a) and S2(b) show the crystal structure at 30 K (LT) with the space group $P\overline{3}$. In the structural refinement, we fixed the Ow–Hw bond length and Hw-Ow-Hw bond angle in H2O molecules at 0.95 Å and 104.5°104.5°104.5  °, respectively. Figure S2(c) shows the $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot resulting from the structural analysis, and Fig. S2(d) displays only the weak superlattice reflections. Here, $F_{\mathrm{c}}$ and $F_{\mathrm{o}}$ indicate the calculated and observed crystal structure factor, respectively. The $|F_{\mathrm{c}}|^{2}$ and $|F_{\mathrm{o}}|^{2}$ of each diffraction reflection show high linearity, indicating good agreement between the experimental and calculated intensities.
@@
If the toroidal arrangement of water molecules is randomly distributed between the nearest hexamers, the structure can be analyzed using a model that assumes disorder of water molecules within the $P\overline{3}1m$ space group. Specifically, this can be represented by a split model for the Ow atoms that constitute the water molecules. We performed structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. The analysis results before and after incorporating the Ow-site split model are shown in Fig. S3(a) and S3(b), respectively. Here, the occupancy of the Ow site is fixed at 1/2. Note that when analyzing a model including hydrogen atoms, the parameters did not converge. Therefore, we present analyses that exclude hydrogen atoms. The split width corresponding to the Ow-Ow distance is 0.161(13) Å. When the Ow-site split model is applied, the $R$ values of the analysis improve; however, it remains significantly worse than that obtained under the assumption of the $P\overline{3}$ space group (Table S II). These results suggest that, in the low-temperature phase, the toroidal order of water molecules is not random but exhibits ferroic arrangement within the plane.
@@
Fig. S 2: Structural analysis assuming $P\overline{3}$ at 30 K. (a) Crystal structure of martyite at 30 K. (b) Arrangement of H2O molecules between layers. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids. $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot of the structural analysis at 30 K showing (c) fundamental and superlattice reflections and (d) only superlattice reflections. (e) Temperature dependence of the Ow–Ow bond distances in the H2O layer. (f) Temperature dependence of the superlattice peak intensity of -4/3 5/3 0. The black dots represent the observed intensity, while the red dots represent the calculations that only consider the Ow displacement from the $P\overline{3}m1$ structure in the HT phase.
Fig. S 3: Structural analysis assuming $P\overline{3}1m$ at 30 K. Results of the structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. Figures (a) and (b) show the analysis results before and after incorporating the Ow-site split model, respectively. Atomic positions are represented as spheres, while atomic displacement parameters (ADPs) are depicted as ellipsoids.
Fig. S 4: O site split model analysis assuming $P\overline{3}$ at 30 K. (a) Results of the Ow split model analysis at 30 K. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids.
Fig. S 5: Framework deformation in the ferrotoroidal phase. Crystal structure of martyite at 30 K. The Zn layer changes little during the phase transition, while in the V2O7 pillar, the O atom between the V atoms displaces by up to 1 % along the $c$-axis direction. The electric dipole moment due to the displacement of the O atoms in the V2O7 pillars interacts with the dipole moment of H2O molecules, resulting in a ferrotoroidal order not only in-plane but also in the stacking direction.
@@ III Energy scale of the intermolecular interactions
Fig. S 6: Dipole-dipole interaction energies between water molecules. (a) In-plane nearest-neighbour interactions. (b) Out-of-plane nearest-neighbour interactions. The numbers in the brackets indicate the degeneracy.
The crystal structure of martyite suggests that the water of crystallization is a quasi-2D system where each water layer is separated by the ZnO4(OH)2 block layers. Here, each H2O molecule can be hydrogen-bonded only with the in-plane nearest neighbors. The second important interaction is the electric dipole-dipole interaction. In this section, we evaluate the dipole-dipole interaction energy scale in martyite. For simplicity, we neglect the electric polarization of the framework of martyite and treat the intermolecular medium as vacuum. The intermolecular distances are fixed at the experimental values determined by the single-crystal XRD at 300 K.
Figure S6(a) shows the dipole-dipole interaction energies of the in-plane nearest-neighbour pair, $E_{\mathrm{iNN}}$. Here, we assume that H2O molecules at A and B sites can take only three orientations as suggested by the XRD at 300 K. The out-of-plane displacement of H2O, which is suggested from the low-temperature XRD, is not taken into account. The interaction energy is estimated as $E_{\mathrm{iNN}}=$ $+1170$ K, $+732$ K, $-146$ K, $-585$ K, from top to bottom molecular geometries. Considering the energy scale of hydrogen bonds from $-2000$ K to $-3000$ K, the bottom geometry is further stabilized. Therefore, water molecules tend to increase the number of hydrogen bonds towards low temperatures. When the hydrogen-bonded H2O hexamers are formed, two types of hexamers appear with CW or CCW electric toroidal dipole moment. The inter-hexamer interaction is determined by the molecular geometry between two hexamers. This energy diagram suggests that the second geometry from the bottom ($-146$ K) is much stabler than the third one ($+732$ K). Indeed, the XRD results suggest that ferrotoroidal order is realized in the LT phase, where the inter-hexamer molecular geometry is of this type.
Second, we estimate the dipole-dipole interaction energy of the out-of-plane nearest-neighbour pair, $E_{\mathrm{oNN}}$. Similarly, we assume that H2O molecules can take only three orientations. The interaction energy is estimated as $E_{\mathrm{oNN}}=$ $+67$ K (top) and $-34$ K (bottom). This energy scale is much smaller than the in-plane one, with the typical ratio of 50 K/2500 K = 2 %, including the hydrogen bonds. Therefore, as indicated from the crystal structure, martyite can be considered as a quasi-2D system. We should note, however, that the out-of-plane interactions can be mediated via the 3D framework deformation as Fig. S5. Nevertheless, the agreement between the experiments and the MD simulations on the monolayer ice suggests that the expression ”monolayer ice in martyite” is not just an exaggeration, but captures the features of this mineral.
@@ IV Raman scattering
We obtained Raman scattering spectra on a single crystal of martyite with the size of 0.5$\times$0.4$\times$0.1 mm3 using a quasi-backscattering geometry (see Fig. S7). We used a 532 nm single-frequency laser (Cobolt Samba) with linear polarization for the incident light. We analyzed the scattered light using a triple monochromator (JASCO NR-1800) equipped with a linear polarizer and a CCD detector (Princeton Instruments Inc. LN/CCD-1100PB). Temperature dependence above 80 K was conducted using THMS600 (LINKAM SCIENTIFIC) with liquid-N2 under a vacuum environment, and below 150 K using a cryostat with GM refrigerator under He atmosphere. We measured Raman spectra on the as-grown $c$ surface of martyite with parallel and perpendicular polarization geometries of the scattered light to the incident light.
@@
Fig. S 7: Quasi-backscattering geometry of the Raman scattering experiment. The sample is mounted on the aluminium stage with a slit for reducing the scattered light from the aluminium surface. The incident angle is approximately 20°20°20  ° with respect to the $c$ axis of the sample. The polarization of the incident light is parallel to the $ab$ plane of the sample.
Fig. S 8: Raman spectra of martyite. The upper panel shows the Raman scattering spectra at 295 K and 4 K. The schematic vibrational modes of H2O ($v_{1}$, $v_{2}$, and $v_{3}$) are also shown. The asterisks indicate the peaks due to the grease for sample mounting. The lower panel shows the assigned peak energies at each temperature. The red symbols indicate the peaks observed only at lower temperatures.
Fig. S 9: Stretching modes of the Raman spectra. The upper panel shows the enlarged Raman scattering spectra around $v_{1}$ and $v_{3}$ from 285 K to 4 K. The lower panel shows the assigned peak energies (p1–p4) at each temperature. The red and green dashed lines are shown for the guide.
@@
In the LT’ phase (even from higher temperatures around 90 K), new peaks appear at 3310 cm-1 and 3440 cm-1. At least, 4 peaks (3390 cm-1 (p1), 3480 cm-1 (p2), 3310 cm-1 (p3), and 3440 cm-1 (p4)) are observed at 4 K. If we assign the 4 modes to $2A_{g}+2E_{g}$ in the $P\overline{3}$ structure, the new modes must appear as the modes are splitting. However, the p3 is a new peak and does not originate from the p1 peak. Therefore, this result suggests that there are at least two H2O sites, and indicates the symmetry lowering probably due to the deformation of the H2O hexamer below $T_{\mathrm{c2}}$ as suggested from the MD simulation.

### v2

@@ Abstract
In conclusion, we study the quasi-2D water of crystallization in martyite, which shows the hierarchical ordering at low temperatures. At room temperature (HT), water molecules are dynamically disordered due to the geometrical frustration of the honeycomb lattice. Here, the rotational and vibrational motions are allowed in the porous framework. At $T_{\mathrm{c1}}$, these water molecules cooperatively form H2O hexamers with hydrogen bonds, leading to the ferrotoroidal order in the LT phase. Once the H2O hexamer is formed, the rotational motions of H2O are forbidden, while the vibrational motions along the hydrogen bonds are still alive. Finally at $T_{\mathrm{c2}}$, these vibrational motions of H2O freeze with the deformation of the H2O hexamers in the LT’ phase. After the deformation, the orientation of the type-1 hexamer (location of the long hydrogen bond) appears as a new degree of freedom. The long hydrogen bonds favor lying close to each other, leading to the formation of the H2O octadecamers. These successive phase transitions gradually break the original honeycomb-lattice symmetry and lead the system close to the ground state, which is difficult to achieve on the highly degenerate energy landscape. This degeneracy is the feature of the monolayer ice, behind the polymorphism of H2O.
@@ Acknowledgements
We acknowledge K. Homma, S. Morita, N. Kawashima, D. Hamane, M. Akashi, J. Nasu, and S. Hayami for the fruitful discussions. This work was partly supported by JSPS KAKENHI, Grant-in-Aid for Scientific Research (Nos. 22K14010, 23H04861, 24K06944, 24H01644, and 24H01650). The synchrotron radiation experiments were performed at SPring-8 with the approval of the Japan Synchrotron Radiation Research Institute (JASRI) (Proposal Nos. 2023A1687 and 2024A1709). The crystal structure figures have been created by using the visualization software VESTA [45].
Supplemental Material of ”Ferroaxial order of the monolayer ice in martyite”
I Methods
Sample preparation. We grew single- and polycrystals of martyite by hydrothermal synthesis. We put 100 mg of ZnO (99.0+%, Fujifilm Wako) powder, 200 mg of V2O5 (99.0+%, Fujifilm Wako) powder, 200 mg of citric acid, and 50 ml of deionized water in Teflon-lined stainless steel autoclaves, and kept them at 140 C for 10–30 days. Hexagonal transparent single crystals up to the size of 0.5$\times$0.5$\times$0.1 mm3 were obtained.
@@ I Methods
Single-crystal XRD. We performed the x-ray diffraction (XRD) experiments using a single crystal of 80$\times$70$\times$10 $\mu$m3 on BL02B1 at a synchrotron facility SPring-8 in Japan [36]. We employed a He-gas-blowing device for controlling the temperature between 30 and 300 K. The X-ray wavelength was $\lambda=$ 0.30918 Å. We used a two-dimensional detector CdTe PILATUS to record the diffraction pattern. The intensities of Bragg reflections with the interplane distance $d>0.3$were collected by CrysAlisPro program [46]. Intensities of equivalent reflections were averaged and the structural parameters were refined by using Jana2006 [47].
Raman scattering. We obtained Raman scattering spectra on a single crystal of martyite with the size of 0.5$\times$0.4$\times$0.1 mm3 using a quasi-backscattering geometry. We used a 532 nm single-frequency laser (Cobolt Samba) with linear polarization for the incident light. We analyzed the scattered light using a triple monochromator (JASCO NR-1800) equipped with a linear polarizer and a CCD detector (Princeton Instruments Inc. LN/CCD-1100PB). Temperature dependence above 80 K was conducted using THMS600 (LINKAM SCIENTIFIC) with liquid-N2 under a vacuum environment, and below 150 K using a cryostat with GM refrigerator under He atmosphere. We measured Raman spectra on the as-grown $c$ surface of martyite with parallel and perpendicular polarization geometries of the scattered light to the incident light.
MD simulation. We carried out molecular dynamics (MD) simulations of two-dimensional water in martyite using GROMACS 2018 software [37]. The model system consists of $N_{w}$ water molecules confined between two atomic layers of martyite. The atomic positions of martyite other than those of water molecules were fixed during the simulations. The number $N_{w}$ of water molecules was 288 unless otherwise stated. The simulation cell was a rectangular prism under periodic boundary conditions in the two dimensions parallel to the atomic layers. The temperature was controlled by the Nosé-Hoover thermostat and the volume of the system was fixed.
@@ II Single-crystal XRD
The results of the structural analysis of martyite are summarized in Tables S I–S IV and Figures S1–S4. Figures S1(a) and S1(b) show the crystal structure at 30 K (LT) with the space group $P\overline{3}$. In the structural refinement, we fixed the Ow–Hw bond length and Hw-Ow-Hw bond angle in H2O molecules at 0.95 Å and 104.5°104.5°104.5  °, respectively. Figure S1(c) shows the $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot resulting from the structural analysis, and Fig. S1(d) displays only the weak superlattice reflections. Here, $F_{\mathrm{c}}$ and $F_{\mathrm{o}}$ indicate the calculated and observed crystal structure factor, respectively. The $|F_{\mathrm{c}}|^{2}$ and $|F_{\mathrm{o}}|^{2}$ of each diffraction reflection show high linearity, indicating good agreement between the experimental and calculated intensities.
@@
If the toroidal arrangement of water molecules is randomly distributed between the nearest hexamers, the structure can be analyzed using a model that assumes the disorder of water molecules within the $P\overline{3}1m$ space group. Specifically, this can be represented by a split model for the Ow atoms that constitute the water molecules. We performed structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. The analysis results before and after incorporating the Ow-site split model are shown in Fig. S2(a) and S2(b), respectively. Here, the occupancy of the Ow site is fixed at 1/2. Note that when analyzing a model including hydrogen atoms, the parameters did not converge. Therefore, we present analyses that exclude hydrogen atoms. The split width corresponding to the Ow-Ow distance is 0.161(13) Å. When the Ow-site split model is applied, the $R$ values of the analysis improve; however, it remains significantly worse than that obtained under the assumption of the $P\overline{3}$ space group (Table S II). These results suggest that, in the low-temperature phase, the toroidal order of water molecules is not random but exhibits a ferroic arrangement within the plane.
@@
Fig. S 1: Structural analysis assuming $P\overline{3}$ at 30 K. (a) Crystal structure of martyite at 30 K. (b) Arrangement of H2O molecules between layers. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids. $|F_{\mathrm{c}}|^{2}$$|F_{\mathrm{o}}|^{2}$ plot of the structural analysis at 30 K showing (c) fundamental and superlattice reflections and (d) only superlattice reflections. (e) Temperature dependence of the Ow–Ow bond distances in the H2O layer. (f) Temperature dependence of the superlattice peak intensity of -4/3 5/3 0. The black dots represent the observed intensity, while the red dots represent the calculations that only consider the Ow displacement from the $P\overline{3}m1$ structure in the HT phase.
Fig. S 2: Structural analysis assuming $P\overline{3}1m$ at 30 K. Results of the structural analysis at 30 K, assuming the $P\overline{3}1m$ space group. Figures (a) and (b) show the analysis results before and after incorporating the Ow-site split model, respectively. Atomic positions are represented as spheres, while atomic displacement parameters (ADPs) are depicted as ellipsoids.
Fig. S 3: O site split model analysis assuming $P\overline{3}$ at 30 K. (a) Results of the Ow split model analysis at 30 K. A H2O hexamer within the black dashed line shows the atomic displacement parameters, indicated by ellipsoids.
Fig. S 4: Crystal structure of martyite at 30 K. The Zn layer changes little during the phase transition, while in the V2O7 pillar, the O atom between the V atoms displaces by up to 1 % along the $c$-axis direction. The electric dipole moment due to the displacement of the O atoms in the V2O7 pillars interacts with the dipole moment of H2O molecules, resulting in a ferrotoroidal order not only in-plane but also in the stacking direction.
@@ III Energy scale of the intermolecular interactions
Fig. S 5: Dipole-dipole interaction energies between water molecules. (a) In-plane nearest-neighbor interactions. (b) Out-of-plane nearest-neighbor interactions. The numbers in the brackets indicate the degeneracy.
The crystal structure of martyite suggests that the water of crystallization is a quasi-2D system where each water layer is separated by the ZnO4(OH)2 block layers. Here, each H2O molecule can be hydrogen-bonded only with the in-plane nearest neighbors. The second important interaction is the electric dipole-dipole interaction. In this section, we evaluate the dipole-dipole interaction energy scale in martyite. For simplicity, we neglect the electric polarization of the framework of martyite and treat the intermolecular medium as a vacuum. The intermolecular distances are fixed at the experimental values determined by the single-crystal XRD at 300 K.
Figure S5(a) shows the dipole-dipole interaction energies of the in-plane nearest-neighbor pair, $E_{\mathrm{iNN}}$. Here, we assume that H2O molecules at A and B sites can take only three orientations as suggested by the XRD at 300 K. The out-of-plane displacement of H2O, which is suggested from the low-temperature XRD, is not taken into account. The interaction energy is estimated as $E_{\mathrm{iNN}}=$ $+1170$ K, $+732$ K, $-146$ K, $-585$ K, from top to bottom molecular geometries. Considering the energy scale of hydrogen bonds from $-2000$ K to $-3000$ K, the bottom geometry is further stabilized. Therefore, water molecules tend to increase the number of hydrogen bonds towards low temperatures. When the hydrogen-bonded H2O hexamers are formed, two types of hexamers appear with CW or CCW electric toroidal dipole moment. The inter-hexamer interaction is determined by the molecular geometry between two hexamers. This energy diagram suggests that the second geometry from the bottom ($-146$ K) is much stabler than the third one ($+732$ K). Indeed, the XRD results suggest that the ferrotoroidal order of H2O is realized in the LT phase, where the inter-hexamer molecular geometry is of this type.
Second, we estimate the dipole-dipole interaction energy of the out-of-plane nearest-neighbor pair, $E_{\mathrm{oNN}}$. Similarly, we assume that H2O molecules can take only three orientations. The interaction energy is estimated as $E_{\mathrm{oNN}}=$ $+67$ K (top) and $-34$ K (bottom). This energy scale is much smaller than the in-plane one, with the typical ratio of 50 K/2500 K = 2 %, including the hydrogen bonds. Therefore, as indicated by the crystal structure, martyite can be considered a quasi-2D system. We should note, however, that the out-of-plane interactions can be mediated via the 3D framework deformation as Fig. S4. Nevertheless, the agreement between the experiments and the MD simulations on the monolayer ice suggests that the expression ”monolayer ice in martyite” is not just an exaggeration, but captures the features of this mineral.
@@ IV Raman scattering
Fig. S 6: Raman spectra of martyite. The upper panel shows the Raman scattering spectra at 295 K and 4 K. The schematic vibrational modes of H2O ($v_{1}$, $v_{2}$, and $v_{3}$) are also shown. The asterisks indicate the peaks due to the grease for sample mounting. The lower panel shows the assigned peak energies at each temperature. The red symbols indicate the peaks observed only at lower temperatures.
Fig. S 7: Stretching modes of the Raman spectra. The upper panel shows the enlarged Raman scattering spectra around $v_{1}$ and $v_{3}$ from 285 K to 4 K. The lower panel shows the assigned peak energies (p1–p4) at each temperature. The red and green dashed lines are shown for the guide.
@@
In the LT’ phase (even from higher temperatures around 90 K), new peaks appear at 3310 cm-1 and 3440 cm-1. At least, 4 peaks [3390 cm-1 (p1), 3480 cm-1 (p2), 3310 cm-1 (p3), and 3440 cm-1 (p4)] are observed at 4 K. If we assign the 4 modes to $2A_{g}+2E_{g}$ in the $P\overline{3}$ structure, the new modes must appear as the modes are splitting. However, the p3 is a new peak and does not originate from the p1 peak. Therefore, this result suggests that there are at least two H2O sites, and indicates the symmetry lowering probably due to the deformation of the H2O hexamer below $T_{\mathrm{c2}}$ as suggested from the MD simulation.